OpenStudy (nuccioreggie):

A theater company charges $5.50 per ticket for people of all age groups. If c represents the total number of children who bought tickets and a represents the total number of adults who bought tickets, which expression best represents the total money earned by the company from the sale of tickets? 5.5(c + a) 5.5c + a c + 5.5a 11(c + a)

OpenStudy (nuccioreggie):

@TwiztTiez @alphadxg

OpenStudy (twizttiez):

Hmmm

OpenStudy (nuccioreggie):

I think its A not sure though

OpenStudy (nuccioreggie):

@TwiztTiez

OpenStudy (twizttiez):

Well lemme think and if I can't come to a conclusion then I'll tag someone ^-^

OpenStudy (alphadxg):

@nuccioreggie Think about what the question is asking for, it is asking for total amount of tickets sold, so it wants to know how many total tickets are sold? which means there can only be one solution

OpenStudy (alphadxg):

@nuccioreggie So in order to get the total amount of tickets, we need to combine C and A, because C = children and A = Adult, either way we need to know the total, so total of C and A = C+A, right?

OpenStudy (alphadxg):

@nuccioreggie So the equation C + A = Total amount of tickets sold, now the question wants to know how much MONEY was made. So each ticket was $5.50. So which equation makes sense the most? Remember, C+A = Total tickets sold, now we need to figure out how many tickets X $$ = Money made

OpenStudy (nuccioreggie):

So is the answer c or a

OpenStudy (alphadxg):

@nuccioreggie Right, but to make it easier, there is no condition set for A adults, so 5.5a doesn't really make any sense, because all the question wants us to know is A + C, and then how much money was made So Adults + Children = Total tickets, now they want to know, "How much money was made from these tickets". To do this simply Multiply by the number of each ticket ($5.50) So (A + C) x $5.50 = total money made

OpenStudy (alphadxg):

@nuccioreggie @TwiztTiez Hopefully that makes sense.

OpenStudy (twizttiez):

Oh so I get to learn as well ^-^

OpenStudy (twizttiez):

Yay!

OpenStudy (nuccioreggie):

Yeup it do but I got the an we I jus want to make sure I got it right ans my answer I got was A

OpenStudy (alphadxg):

niceeeee

OpenStudy (nuccioreggie):

Is A correct

OpenStudy (alphadxg):

@nuccioreggie Yes you are right

OpenStudy (nuccioreggie):

Thanks